OpenStudy (anonymous):

if sinx = -3/5 and x is in quadrant 3, then tan2x= ___ a. 24/7 b. 12/7 c. -24/7 d-12/7

OpenStudy (anonymous):

\[\sin x = (\frac{ opp }{ hyp })\] This tells you that your sides are of length: 3, 4, and 5. \[\cos(x) = -4/5\]\[\tan(x) = \frac{ \sin(x) }{ \cos(x) } =(-3/5)/(-4/5) = 3/4\] Tan 2x is a double-angle formula. \[\tan(2x) = 2\tan(x)/(1-\tan ^{2}(x))\]\[(2(3/4))/ (1-(3/4)^{2}) = (1.5/0.4375) = 3.43\]

OpenStudy (anonymous):

Otherwise known as (24/7) or part A.

OpenStudy (anonymous):

ty for giving me the wrong answer

OpenStudy (anonymous):

That doesn't make sense. It should be right.

OpenStudy (anonymous):

well..its not

OpenStudy (anonymous):

Are you sure you didn't mean x is in quadrant 2?

OpenStudy (anonymous):

Because if x is in quadrant 2, then the answer changes.

OpenStudy (anonymous):

im sure :/